Johnny has $100 now and saves $10 every month. His brother Jimmy has no money
currently and is going to save $15 per month. In how many months will Jimmy have
as much money as Johnny?

Answer by mananth(16946)

johnny 100 + 10 /month
Jimmy 15 /month
let them have same amount after x months
...
100+10x=15x
15x-10x=100
5x=100
x= 20 months
...
m.ananth@hotmail.ca
..
check
15*20=300
20*10+100=300